OpenStudy (anonymous):

Joan has 2 similar rectangular-shaped pieces that she is using for her quilt. One hias width 4 inches and length 6 inches and other has a width of 6 inches. What is the perimeter of the larger rectancular piece ?

OpenStudy (anonymous):

doesnt

OpenStudy (anonymous):

Do you have options or answers ?

OpenStudy (anonymous):

32 inchs ?

OpenStudy (espex):

\[\frac{4}{6}=\frac{6}{x} \rightarrow 4x=36 \rightarrow x=9\] so (2)(6)+(2)(9)=12+18=30 inches